### v1

@@ 3.2 Emotions
Changes to the feed algorithm caused in-feed, but not post-experiment, changes to participants’ negative emotions (Fig. 3). Reduced Exposure caused a change of -5.05 points (95% CIs: [-7.68, -2.42]; $P_{adj}$=0.002) for “angry” and -3.68 points (95% CIs: [-6.42, -0.94]; $P_{adj}$=0.024) for “sad.” Likewise, in the Increased Exposure, participants report a symmetric increase in negative emotions: 5.13 points (95% CIs: [2.05, 8.22]; $P_{adj}$=0.013) for “angry” and 4.38 points (95% CIs: [1.00, 7.76]; $P_{adj}$=0.036) for “sad.” Neither intervention had an effect on the positive emotions “calm” and “excited.” Post-experiment emotion effects were not significant, suggesting that the effects were likely short-lived.
3.3 Engagement
We tested whether our interventions impacted platform engagement (Sec. E.4). Neither experiment observed imbalances in attrition rates across conditions. Moreover, in both experiments, the treatment had no measurable impact on daily sessions—Increased Exposure: 1.65 control vs. 1.53 treatment, n.s.; Reduced Exposure: 1.53 vs. 1.49, n.s. Likewise, there was no significant effect on reshare rates—Increased Exposure: 0.52% vs. 1.07%, n.s.; decreased exposure: 0.74% vs. 0.46%, n.s. In addition, there was no effect on favorite rates—Increased Exposure: 4.48% vs. 5.46%, n.s; decreased exposure: 4.96% vs. 4.15%, n.s. We also observed that the differences between treatment and control groups in terms of the number of new posts, replies, and time spent on X during the experimental period were not statistically significant. These null results persisted when controlling for baseline behavior. While nominally consistently in the expected direction, none of the statistical tests were significant. It is possible that an even larger study could observe differences; however, our data suggests that the magnitude of any effect would be small.
In an exploratory analysis, we did observe that political content had higher engagement than non-political content, confirming the findings of previous work [49]. When focusing only on political content, we found that AAPA posts do not get more engagement than moderate political posts (Sec. E.4).
4 Discussion
@@
In this study, we conducted a field experiment to investigate the effects of feed ranking algorithms that decrease or increase exposure to contact that polarizes. Our findings demonstrate that repeated exposure to AAPA posts, particularly during heightened political tensions such as months before the US election, significantly influences users’ affective polarization and emotional responses. A feed ranking algorithm that controls exposure to such content produces measurable changes, equivalent to mitigating approximately three years of affective polarization, even after just one week of treatment. Importantly, our results indicate that the intervention does not disproportionately impact any specific political leaning or demographic group, suggesting that the approach is bipartisan. The success of this method shows that it can be integrated into social media AIs to mitigate harmful societal consequences. We also present evidence that this impact can likely be achieved with a limited effect on business metrics like engagement.
Limitations of our study include a sample of partisan US users restricted to the web version of X during a specific period, which may affect the generalizability of the findings [50, 51]. Our population is limited to X users with at least 5% of political content (73.4% of the screened), but we intentionally selected this group to ensure the interventions can make a difference. Additionally, the experiment examines short-term effects—immediate in-feed responses and outcomes measured the day after the experiment—but does not address whether these effects persist over time. However, in scenarios where this ranking approach is integrated into the feed algorithm, the intervention would be continuously applied. Our reliance on in-feed surveys to assess immediate responses to stimuli follows established practices of ecological momentary assessment (EMA) [45]. To mitigate the risk of demand characteristics associated with the close proximity of the assessment, we followed best practices in designing the experiment [52], providing only essential information about the study’s purpose. Post-survey responses to an open-ended question about their experience during the study suggest that these precautions were effective, as participants did not report any impact on their experience attributable to their assigned condition. Describing the causal relation between social media and polarization as a unified phenomenon, of course, remains challenging: platforms are complex socio-technical systems, and small variations in design or exogenous factors can lead to different outcomes [53]. Future research should explore whether these effects persist over longer periods or in less politically charged contexts and examine the efficacy of feed reranking across different platforms and cultural settings.
@@
As political polarization and societal division become increasingly linked to social media activity, our findings provide a potential pathway for platforms to address these challenges through algorithmic interventions. By embedding pro-democracy attitudes into AI-driven feeds [37, 39], we offer a concrete solution that could contribute to healthier democratic discourse, enhance social trust, and help mitigate the risks of deepening polarization.
Appendix A Interventions
A.1 Posts expressing Antidemocratic Attitudes and Partisan Animosity

### v2

@@ Results
Changes to the feed algorithm caused in-feed, but not post-experiment, changes to participants’ negative emotions (Fig. 3). Reduced Exposure caused a change of -5.05 points (95% CIs: [7.68-7.682.42-2.42]; $P_{adj}$=0.002) for “angry” and -3.68 points (95% CIs: [-6.42, -0.94]; $P_{adj}$=0.024) for “sad.” Likewise, in the Increased Exposure, participants reported a symmetric increase in negative emotions: 5.13 points (95% CIs: [2.05, 8.22]; $P_{adj}$=0.013) for “angry” and 4.38 points (95% CIs: [1.001.007.767.76]; $P_{adj}$=0.036) for “sad.” Neither intervention had an effect on the positive emotions “calm” and “excited.” Post-experiment emotion effects were not significant, suggesting that the effects were likely short-lived.
Engagement
We tested whether our interventions impacted platform engagement (Appendix S9). Neither experiment observed differential attrition rates across conditions.
In the Reduced Exposure experiment, participants in the treatment group had a similar number of daily sessions (estimate: -0.06; 95% CI: [-0.16, 0.04]), but spent 4.71-4.71 fewer minutes per day (95% CI: [-8.22, -1.20]) compared to the control group. Over the 7-day intervention period, they viewed -169.52 fewer posts (95% CI: [-325.48, -13.56]), liked -24.06 fewer posts (95% CI: [39.53-39.538.59-8.59]), but reposted a similar number of posts (estimate: -4.18; 95% CI: [-9.89, 1.53]). On the other hand, they engaged with the posts they viewed at higher rates, with a 0.61% higher like rate (95% CI: [0.11%, 1.12%]) and a 0.058% higher repost rate (95% CI: [0.009%, 0.11%]). While the per-view engagement rates increased, these findings suggest that downranking AAPA content comes at the cost of slightly reduced engagement overall.
In the Increased Exposure experiment, we did not find any significant differences in engagement between the treatment and control groups. Participants in the treatment group had a similar number of daily sessions (estimate: -0.04; 95% CI: [-0.16, 0.08]) and spent a comparable amount of time on the platform (estimate: -1.76 minutes per day; 95% CI: [-6.17, 2.66]). They also viewed (estimate: 132.3-132.3; 95% CI: [-292.5, 27.94]), liked (estimate: 10.25; 95% CI: [-5.23, 25.74]), and reposted (estimate: 3.38; 95% CI: [-0.06, 6.82]) a similar number of posts, and liked and reposted the posts they viewed at similar rates (effect on like rate: 0.22%, 95% CI: [-0.57%, 1.01%]); effect on repost rate: -0.052%; 95% CI: [-0.17%, 0.066%]).
Discussion
This article demonstrates a deployable framework for conducting reranking field experiments on social media feeds without requiring direct involvement from the platforms. We developed a browser extension infrastructure that enables real-time manipulation of what consenting users see in their feed, using LLM-based classifiers to dynamically assess and rerank posts based on a custom ranking function. This approach opens new avenues for auditing, evaluating, and prototyping alternative feed-ranking strategies in real-world settings.
@@ Discussion
Our specific experimental findings demonstrate that repeated exposure to AAPA posts, particularly during heightened political tensions such as the months leading up to the 2024 U.S. election, significantly influences users’ affective polarization and emotional responses. A feed-ranking algorithm that controls exposure to such content produces measurable changes, equivalent to approximately three years of affective polarization, over the period it was deployed. Importantly, our results indicate that the intervention does not disproportionately impact any specific political leaning or demographic group, suggesting that the approach is bipartisan. The success of this method shows that it can be integrated into social media AIs to mitigate harmful personal and societal consequences. At the same time, our engagement analyses indicate a practical trade-off: interventions that downrank AAPA may reduce short-term engagement volume, posing challenges for engagement-driven business models and supporting the hypothesis that content that provokes strong reactions generates more engagement.
Methodological advances and policy implications
Our work opens the door to conducting large-scale causal inference on social media feed algorithms without requiring “independence by permission”  (?) from platforms. While prior work made important advances in randomized controlled trials on the Facebook and Instagram algorithms (?), these experiments could only proceed with negotiated support from Meta’s executives—an arrangement that may limit the scope and independence of scientific work. Our approach enables scientists to test a much broader set of interventions beyond those that platforms might agree to test, fostering a more robust ecosystem for independent social media research.
Limitations and future directions
Our study focused on a sample of partisan U.S. users restricted to the web version of X during the months leading up to a U.S. presidential election. This context, marked by a high volume of political content and emotional engagement, enabled us to test interventions effectively, but may limit the generalizability of the findings  (?, ?) beyond a period of heightened political activity. Our population is limited to X users with at least 5% of political content, but we intentionally selected this group to ensure the interventions were viable. Additionally, the experiment examines short-term effects—immediate in-feed responses and post-survey responses measured soon after the experiment—but does not address whether these effects persist over time. In additional analysis examining the variation of the treatment effects by experiment day and the time gap between the end of the experiment and when the participants started the post-survey, we did not find any evidence for immediate decay of the observed effects (Appendix S11.5). We also note that in scenarios where this ranking approach is integrated into the feed algorithm, the intervention would be continuously applied, and any effects may dissipate once it is turned off. Other technical limitations of our approach include the difficulty in performing cluster-randomized experiments testing for social influence  (?), and the inability to collect the full feed inventory for a given user.
Our reliance on in-feed surveys to assess immediate responses to stimuli follows established practices of ecological momentary assessment  (?). To mitigate the risk of demand characteristics associated with the close proximity of the assessment, we followed best practices in designing the experiment  (?), providing only essential information about the study’s purpose. Post-survey responses to an open-ended question about participants’ experiences during the study suggest that these precautions were effective, as they did not report any impact on their experience attributable to their assigned condition (Appendix S11.4).
Describing the causal relation between social media and polarization as a unified phenomenon remains challenging: platforms are complex socio-technical systems, and small variations in design or exogenous factors can lead to different outcomes  (?). Future longitudinal research should investigate the impact of sustaining the intervention over longer periods of time—especially on long-term metrics such as user retention  (?) and news knowledge—and evaluate the efficacy of feed reranking across different platforms and cultural settings. For example, open questions remain about how our approach can be extended to the broader online information ecosystem and if our findings hold true in different political systems with unique cultural contexts or party structures. In addition, future research can move beyond the initial eight factors related to antidemocratic attitudes and operationalize new factors grounded in other social science theories  (?), and measure how this content can impact other experiences online, such as mental health, trust in institutions, civic engagement, or the quality of democratic discourse. Using a similar design as our study, these future investigations can also examine different time horizons, from immediate reactions captured through in-feed surveys to long-term impacts assessed through follow-up surveys conducted days or even months later.
@@
As political polarization and societal division become increasingly linked to social media activity, our findings provide a potential pathway for platforms to address these challenges through algorithmic interventions. Together, these interventions may result in algorithms that not only reduce partisan animosity but also promote greater social trust and healthier democratic discourse across party lines  (?, ?).
Materials and Methods Summary
Participant recruitment.
We recruited participants through two online platforms, CloudResearch and Bovitz, targeting U.S. residents over 18 years old who self-identified as either Republican or Democrat and were active users of X (Section S1.1). Qualified individuals were invited to complete a screening task, which included installing a browser extension that analyzed their X feed. To ensure the interventions could have a meaningful impact on participants’ feeds, only those with at least 5% of posts related to politics or social issues were invited to participate. Figure S1 summarizes the recruitment funnel, including the number of individuals at each stage of the process. Participants were not instructed to use X in any particular way, but they received daily reminders if they had not used the platform that day.
Surveys.
Participants were required to complete a pre-experiment survey that included questions about their demographics, personal values, affective experiences on social media, and political attitudes (Section S1.2). At the end of the study, they were asked to complete a post-experiment survey with a similar structure, but without questions on demographics or personal values. The surveys took about 20 and 10 minutes to complete, respectively. The section of the survey measuring political attitudes included a question related to our primary outcome, assessing participants’ feelings toward out-partisans, while the section on affective experiences on social media included questions related to our secondary outcomes, measuring how often participants felt angry, sad, excited, and calm.
Throughout the experiment, we also inserted short surveys into participants’ feeds to measure their feelings in the moment. Each in-feed survey included either one question about participants’ feelings toward members of the out-party or two questions asking how much they felt one positive emotion (excited or calm) and one negative emotion (angry or sad), randomly selected. To minimize survey fatigue, once a participant answered a question, they were not shown the same type of question for the next ten minutes and were subsequently less likely to receive that question again.
Feed reranking implementation.
To administer the interventions, we developed a browser extension that intercepted the participants’ original X feeds, sent the posts to a remote backend for real-time analysis, and displayed the reranked feeds (Section S3.2). The backend used a fast, custom-built classifier to detect political posts (Section S4.3), and OpenAI’s GPT-3.5 Turbo model to score the political posts on the eight AAPA dimensions (Section S4.4). The reranking process took approximately three seconds and was consistent across all experimental conditions. The extension also administered in-feed surveys and logged participants’ engagement with the feed, including which posts they viewed, liked, and reposted.
Experimental design.
We used Bernoulli randomization to assign participants to either the Reduced or Increased Exposure experiment, and subsequently to treatment and control groups with equal probability (Section S1.3). Due to differences in how the two interventions were operationalized, each experiment had its own control group. To determine the sample size, we conducted a power analysis based on data from a pilot study with 70 participants that followed the same experimental design and analysis procedure. We chose target quotas that enabled us to detect effects of at least two degrees on the feeling thermometer scale for affective polarization with 95% power. We used randomization inference to test for imbalances in participants’ characteristics, differential attrition rates, and differential attrition patterns between the treatment and control groups in each experiment. We found no evidence of covariate imbalance or differential attrition.
Analysis.
As preregistered, to estimate the effects of the interventions on the post-experiment outcomes, we used ordinary least squares, regressing the outcome variable on the treatment assignment indicator, the participant’s response to the same question on the pre-survey, and a recruitment platform indicator (Section S1.4). To estimate the effects on the outcomes measured with in-feed surveys, we used linear mixed-effects models to account for the multiple observations per participant. We regressed the outcome on the treatment assignment, the participant’s average response on the same question during the three-day baseline period, and the recruitment platform indicator. To estimate heterogeneous treatment effects, we used the same model specifications but included an interaction between the treatment assignment and the moderator, running one regression per moderator (Section S8). As preregistered, we considered party identification, party strength, and the proportion of political content in the participant’s feed during the baseline period as primary moderators, and gender, race, age, income, and socioeconomic ladder as secondary moderators. To account for testing multiple hypotheses, we applied False Discovery Rate adjustment (Section S6). Finally, to appropriately model the nature of the outcomes, we used quasi-Poisson models to estimate the effects on engagement volume and Beta models to estimate the effects on engagement rates (Section S9).
Acknowledgments
We thank Pablo Barberá, Dean Eckles, Matthew Gentzkow, Desmond Ong, Jennifer Pan, Sandy Pentland, Johan Ugander, Robb Willer, and the anonymous reviewers for helpful feedback and discussions. Michael Bernstein discloses an additional affiliation as a co-founder of Simile AI, Inc., which builds AI human behavioral simulations.
Funding:
This work was supported in part by the NSF under awards IIS-2403433, IIS-2403434, IIS-2403435; BCS-2214203; Swiss NSF under award P500PT-206953, and a Hoffman-Yee grant from the Stanford Institute for Human-Centered Artificial Intelligence.
Author contributions:
Conceptualization: T.P., M.S., C.J., J.T., J.H., M.B.
Methodology: T.P., M.S., C.J., J.T., J.H., M.B.
Software: T.P.
Validation: T.P., M.S.
Formal analysis: T.P., M.S.
Investigation: T.P., M.S.
Resources: T.P., M.S., C.J.
Data curation: T.P., M.S.
Writing – original draft: T.P., M.S., C.J., M.B.
Writing – review & editing: T.P., M.S., C.J., J.H., J.T., M.B.
Supervision: J.T., J.H., M.B.
Project administration: T.P., M.S., C.J., M.B.
Funding acquisition: M.S., C.J., J.H., M.B
Competing interests:
The authors declare that they have no competing interests.
Data and materials availability:
The repository associated with the manuscript, containing aggregated data and code required to reproduce the paper results, is available at: (?)
The pre-registration of the study is available at: (?)
Supplementary Materials for
Reranking partisan animosity in algorithmic social media feeds alters affective polarization
Tiziano Piccardi∗†, Martin Saveski∗†, Chenyan Jia,
Jeffrey Hancock, Jeanne Tsai, Michael Bernstein
Corresponding author. Email: piccardi@jhu.edu (T.P.); msaveski@uw.edu (M.S.)
These authors contributed equally to this work.
This PDF file includes:
Methods and Materials
Ethical Considerations
Supplementary Text
Figs. S1 to S24
Tables S1 to S32
References (59-81)
Supplementary Material
S1 Methods
S1.1 Participant Recruitment
We recruited participants from two online platforms, Bovitz and CloudResearch. Using platform-specific filters, we targeted individuals who met the following criteria: US residents aged 18 or older, self-identifying as either Republican or Democrat, and active users of X. Participants were invited to complete a two-minute screening task to assess their eligibility for the study. The experiment was restricted to desktop users on Google Chrome or Microsoft Edge browsers. Upon accepting the consent form, participants were instructed to install a web extension that launched X in a new browser window, automatically scrolled through their “For you” feed to capture approximately 180 posts, and calculated the proportion of posts related to politics or social issues using our classifier (Appendix S4.3). Only participants with at least 5% of posts related to politics or social issues were invited to join the study, ensuring that participants regularly engaged with political content, a necessary condition for the intervention to have an impact on their feed. Participants were compensated \$0.50 for completing the screening task and \$20 upon finishing the study. Fig. S1 provides a visual summary of the full recruitment process and the number of participants at each step.
Figure S1: Summary of the participants’ recruitment funnel.
On Bovitz, 4,293 participants initiated the screening task, and 1,753 successfully installed the extension and accessed X. The observed drop in participation was primarily due to the use of unsupported devices or browsers (e.g., mobile). Participants for whom we collected fewer than 100 posts due to technical issues (e.g., conflicts with other browser extensions) or who deliberately interrupted the screening were excluded from the study, leaving 1,688 eligible participants. After applying the 5% political content filter, 1,373 participants were invited to continue with the study, of which 1,155 completed the pre-experiment survey and were successfully enrolled. Of these, 720 were randomly assigned to the Reduced Exposure experiment (368 in the treatment and 352 in the control group) and 435 to the Increased Exposure experiment (223 in the treatment and 212 in the control group), as our pilot study revealed that the first group required more participants to detect the effect of the intervention  (?).
On CloudResearch, 1,333 participants began the screening, and 1,036 successfully installed the extension and accessed X. The lower dropout rate compared to Bovitz is attributed to the platform’s use of more specific filters, such as restricting participation to desktop users and highlighting the need for software installation, as well as advanced filters targeting active X users. After accounting for successful screening completion (998 participants) and applying the 5% political content criterion, 604 participants were invited to the study. Among these, 507 participants joined, with 275 in the Reduced Exposure experiment (133 in the treatment and 142 in the control group) and 232 in the Increased Exposure experiment (117 in the treatment and 115 in the control group).
Fig. S15 shows the distribution of political posts among participants before the filtering (mean = 17.1%, median = 12.1%), illustrating that the 5% political content filter excluded 26.5% of the initial participant pool. Participants were invited to engage daily and received reminders via email. After ten days, participants with minimal participation, i.e., those for whom we recorded fewer than ten feed load events, were excluded from the study and did not receive the post-experiment survey. The final completion rates were 69.1% (798) for Bovitz and 90.3% (458) for CloudResearch.
Of the participants who completed the study, 602 (47.9%) identified as female, 631 (50.2%) as male, and 23 (1.9%) as other gender identities. In terms of political affiliation, 831 (66.1%) were Democrats, and 425 (33.9%) were Republicans. Fig. S14 summarizes additional demographics such as race, education level, income, and perceived socioeconomic status.
S1.2 Surveys
Participants were required to complete a pre-experiment survey, which lasted approximately 20 minutes. This survey collected demographic information, personal values  (?, ?), affective experiences on social media, and political attitudes such as feelings toward members of the opposing political party and the remaining seven measures of antidemocratic attitudes. The political feeling question was asked first, whereas the order of the other questions was randomized. To ensure contextual relevance, survey questions were adapted based on participants’ political leanings. We followed the structure of the survey used in the study that introduced the eight AAPA variables  (?). At the end of the study, participants completed a post-experiment survey, which took approximately 10 minutes. This survey asked about political attitudes and affective experiences, excluding demographics and personal values. Political attitudes were assessed using a feeling thermometer on a scale from “Very cold or unfavorable feeling” (0) to “Very warm or favorable feeling” (100). To evaluate emotions, participants rated how frequently they felt 15 different emotions on a scale from “Never” (0) to “All the time” (5).
For the in-feed response, the survey for the political feeling assessment asked “At the moment, how do you feel about [out-party]?” where out-party is replaced with “Republicans” for Democratic participants and vice versa. For emotional responses, the survey asks, “How much do you feel” for one negative (sad or angry) and one positive (excited or calm) emotion. These in-feed surveys were designed to resemble posts with a yellow background. Both political feelings and emotions were measured on a scale from 0 to 100, and participants clicked on a slider to indicate their response. The numeric value was not displayed to prevent anchoring effects, and the slider was initially shown empty with no default setting. For political feelings, the slider was labeled with “Very cold or unfavorable feeling” (0), “No feeling at all” (50), and “Very warm or favorable feeling” (100). For emotions, the labels were “None at all” (0), “A little” (25), “Moderately” (50), “A lot” (75), and “Extremely” (100). To avoid survey fatigue, the survey was not shown with every post. Each day, participants had an equal probability (50%) of receiving either a political feeling or emotion question. Once a survey was answered, the same type of question would not appear for the next 10 minutes, and the likelihood of receiving another survey that day would be halved. When selected, the survey was inserted into the feed according to the logic described in the following sections.
S1.3 Experimental Design
Randomization. We used Bernoulli randomization to allocate participants to the experiments with either Reduced or Increased Exposure to AAPA content until we reached the respective target quotas from our power analyses  (?) and to a treatment arm (treatment or control) with an equal probability. We opted for a Bernoulli randomization since we did not have access to the complete participant pool and their characteristics in advance.
Control groups. We ran the Reduced and Increased Exposure experiments as separate, parallel studies, each with its own control group. This was necessary because the interventions were operationalized differently. In the Increased Exposure experiment, we could uprank AAPA content at any point in the feed, allowing for consistent placement of in-feed surveys. In the Reduced Exposure experiment, by contrast, we could only intervene when AAPA content was already present, which limited both the intervention and the placement of surveys. To ensure internal validity, we designed each control group to mirror the structural features of its corresponding treatment in terms of survey timing and content availability.
Power analyses. To inform the sample size for our experiments, we ran a power analysis based on data from a pilot study with 70 participants that followed the same experimental design and analysis procedure. Due to the complexity of the mixed-effects models, we used simulations to conduct the power analyses (using the simR package) with models that correspond to those described in Section S1.4. We optimized for power in detecting effects on our primary dependent variables, i.e., the in-feed and post-experiment effects on affective polarization.
Due to the uncertainty in the recruitment process, we preregistered two sample size targets that, based on a pilot study, would enable us to detect effects of 2 and 2.5 degrees on the feeling thermometer scale with 95% power. We were able to reach and slightly exceed the required sample size to detect effects of at least 2 degrees in both the Reduced and Increased Exposure experiments.
Covariate balance. We used randomization inference to test whether the imbalance in the background characteristics of the participants assigned to treatment vs. control in each of the two experiments was larger than we would expect from chance  (?). We regressed the treatment indicator on all background characteristics and computed a heteroskedasticity-robust Wald statistic for the hypothesis that all the coefficients are equal to zero. Then, we sampled 10,000 assignments using the same randomization procedure, performed the same regression, and compared the Wald statistics to the observed.
We considered all questions in the pre-experiment survey as covariates, including demographics (gender, race/ethnicity, age, education, SES ladder, and income), party affiliation (party ID, party ID strength, party importance), affective polarization (in- and out-party feeling thermometer), antidemocratic attitudes (support for undemocratic practices, support for undemocratic candidates, support for partisan violence, biased evaluation of politicized facts, social distance, social distrust, opposition to bipartisan cooperation), emotions (enthusiastic, happy, still, lonely, sad, nervous, satisfied, calm, relaxed, tired, fearful, aroused, excited, bored, and angry), and recruitment platform. We coarsened the education, income, and SES ladder covariates to avoid small categories. For the sample of participants that completed enough of the in-feed surveys, we also included their average answer to the affective polarization in-feed questions during the three-day baseline period.
We did not find evidence of covariate imbalance in the Reduced Exposure ($p=0.59$) or the Increased Exposure experiment ($p=0.18$). For reference, the distributions of the demographic covariates and univariate comparisons can be found in Tables S4 and S5.
Attrition. We also used randomization inference to test for differential attrition rates and differential attrition patterns.
In the Reduced Exposure experiment, the attrition rate was 36.1% in the treatment group and 38.9% in the control group. In the Increased Exposure experiment, the attrition rate was 30.0% in the treatment group and 29.7% in the control group.
To test whether the difference in attrition rates is larger than we would expect by chance, we computed the t-score of a two-tailed unequal-variances t-test assessing whether the treatment affects the attrition rate, and compared it against the empirical distribution under 10,000 random reassignments of the treatment  (?). We found no evidence of differential attrition rates in either of the two experiments: Reduced Exposure, $p=0.38$; and Increased Exposure, $p=0.92$.
To test for differential attrition patterns, we regressed an attrition indicator on treatment, baseline covariates, and treatment-covariate interactions, and performed a heteroskedasticity-robust F-test of the hypotheses that all interaction coefficients are equal to zero  (?). We considered the same set of covariates as those described in the covariate balance test: demographics (gender, race/ethnicity, age, education, SES ladder, and income), party affiliation (party ID, party ID strength, party importance), affective polarization (in- and out-party feeling thermometer), antidemocratic attitudes (support for undemocratic practices, support for undemocratic candidates, support for partisan violence, biased evaluation of politicized facts, social distance, social distrust, opposition to bipartisan cooperation), emotions (enthusiastic, happy, still, lonely, sad, nervous, satisfied, calm, relaxed, tired, fearful, aroused, excited, bored, and angry), and recruitment platform. Then, we compared the observed F-statistics with the empirical distribution under 10,000 random reassignments of the treatment. We did not find any evidence of differential attrition patterns in either of the two experiments: Reduced Exposure, $p=0.34$; and Increased Exposure, $p=0.72$.
S1.4 Analyses
All analyses follow our preregistration plan filed at the Open Science Foundation  (?). The reported results focus on the participants who completed the post-survey and at least one in-feed survey for the corresponding outcome. We observe similar results when we consider only those who completed the post-survey when analyzing the post-experiment effect and only those who completed at least one in-feed survey when analyzing the in-feed effect (Sections S11.1 and S11.2).
Post-experiment effects. To estimate the average treatment effects of the intervention on the post-experiment versions of outcomes (i.e., measured via pre- and post-experiment survey), we used OLS regressing the outcome variable on the treatment assignment indicator, the participant’s response to the same question on the pre-experiment survey, and a variable indicating the platform they were recruited from (“platform indicator”). We also preregistered that we would consider including other control covariates selected via LASSO regression on the outcomes, but found that only the pre-experiment survey answers were selected. The regression tables for the affective polarization and emotions outcomes are reported in Sections S11.1 and S11.2, respectively.
In-feed effects. To estimate the treatment effects on the outcomes measured using in-feed surveys, we used linear mixed-effects models that account for the fact that there are multiple observations for each participant. We regressed the outcome on the treatment indicator, the participants’ average answer to the same question during the three-day baseline period, and the platform indicator. In cases where we do not have any responses from the participants in the baseline period, we used a linear model to impute their average based on their response to the equivalent pre-survey question. The model was trained to predict the participant’s baseline average given their pre-survey response and fitted on data from all other participants in the corresponding experiment. Participants without answers in the experimental period were excluded from the analysis. The regression tables for the affective polarization and emotions outcomes are reported in Sections S11.1 and S11.2, respectively.
Engagement analysis. We analyzed traditional engagement metrics by focusing on events captured by the web extension. The extension logs each post that enters the browser viewport and remains visible for at least one second, as well as any user interactions such as likes or reposts. Our analysis investigates the return rate and time spent on the platform, key metrics for social media platforms  (?), along with volume and rate of likes and reposts, which play a significant role in how X’s feed algorithm ranks posts, as documented in the open-source code released by X  (?).
S2 Ethical Considerations
To minimize the risks, we have complied with all relevant ethical regulations. All procedures were approved by our university’s Institutional Review Board. All participants were provided with informed consent before installing the browser extension and participating in the study. Participants were given the option to withdraw their consent or discontinue participation at any time without penalty. At the end of the experiment, participants received financial compensation through the recruitment platform and a debriefing statement via email.
For any social media field experiment, we must consider the ethical implications given the high ecological validity of such work  (?). Just like other algorithmic feed ranking methods (e.g.,  (?)), our work may also inherently pose risks to participants by altering the ranking of content on their news feeds. We balance this risk against an acknowledgment that today’s social media platforms (e.g., X) already intervene in people’s feeds and take control over users’ information stream  (?). While this work shares limitations with existing algorithmic ranking methods in that participants may cede control over their feeds to a certain extent, our work does not introduce substantial additional risks. Our down-ranking condition reduced exposure to AAPA posts, while the up-ranking condition inserted AAPA posts that already existed on X and up-ranked an order of magnitude fewer posts. Our aim was to mitigate risks while still carefully and causally identifying whether, and how, changes to feed ranking affect polarization.
S3 Interventions
S3.1 Posts Expressing Antidemocratic Attitudes and Partisan Animosity